# Semantic Mapping in Indoor Embodied AI – A Survey on Advances, Challenges, and Future Directions

### v2

@@ 2.3.2 Modular pipeline
Another line of work explores how to breakdown a complex navigation task into a set of basic skills that the agent needs to acquire. Such skills can then be learned independently of each other so that they can be leveraged across various tasks without the need to be retrained from scratch. While some use two modules (Bansal et al., 2020), one for high-level decision making and another for low-level planner, others (Chaplot et al., 2019; 2020a; Gervet et al., 2023; Raychaudhuri et al., 2023) further breakdown into sub-modules, such as visual encoder, mapper and exploration in addition to the low-level planner. Fig. 4 shows the difference between the end-to-end approach and the module approach, and we elaborate on the modules and popular design choices below.
Visual Encoder. This module encodes agent observations to produce semantic visual features and predictions at every time step. Prior works have used visual features from pretrained backbones such as ResNet He et al. (2016) or ViT (Dosovitskiy et al., 2020), and often leveraging object detectors MaskRCNN He et al. (2017a) or FasterRCNN Ren et al. (2015). As we will see in Sec. 5.2.2, with the development of large pretrained vision-language models and open-vocabulary detectors, pretrained models such as CLIP (Radford et al., 2021), LSeg (Li et al., 2022), DINO (Caron et al., 2021; Oquab et al., 2023), and others are increasingly popular as the basis for building large to open-vocabulary maps. The visual encoder used will determine the information captured in the features, and whether there are detected object instance bounding boxes or segmentations for integration into the mapping modules.
Mapper. The mapper is responsible for building a semantic map of the environment from the encoded image features and agent pose. To build a global map over time, the mapper typically aggregates the current map with the map from previous step (see Sec. 3.4 for details). In this paper, we survey how recent methods structure the map (Sec. 4) and what information can be encoded in it (Sec. 5). Tab. 2 summarizes the type of information stored in the map by various methods. This can be occupancy information, explored area or semantic labels of the detected objects.
@@
In Tab. 2, we compare common modular approaches and look at how they leverage various heuristics-based or learned approach for each of the modules. The advantages of a modular pipeline include its ability to leverage pretrained models from other tasks (Gervet et al., 2023; Raychaudhuri et al., 2023) and its ability to transfer from simulation to real-world robots better (Gervet et al., 2023).
@@ 3 Semantic map
In this section, we introduce semantic maps (Sec. 3.1), and provide an overview of their structural representations (Sec. 3.2), semantic encodings (Sec. 3.3), and the process of building them (Sec. 3.4).
@@ 3.1 What are semantic maps?
While maps, capturing the geometry of the space, can help an agent avoid obstacles, they are not sufficient for the demands of more complex reasoning tasks. An enhanced map that goes beyond geometry to capture meaning and context in its environment aligns with how humans perceive and navigate its surroundings. These are semantic maps, which provide a richer and more nuanced understanding about the objects and places in the environment and are indispensable for performing complex tasks such as navigating to a specific room (kitchen) (Narasimhan et al., 2020), rearranging objects (Trabucco et al., 2022) or performing an action on a specific object (sitting on a couch) (Peng et al., 2023). A physical or a virtual agent perceives the environment through sensors (camera, LiDAR, etc.), uses cognition to classify the objects and regions it perceives (Ren et al., 2015; He et al., 2017a; Liu et al., 2023b; Kirillov et al., 2023; Zhang et al., 2023b), stores them in a structured semantic map which can be queried for efficient reasoning.
@@ 3.2 What is the structure of this map?
A semantic map can be structured as a spatial grid map, topological map, dense geometric or a hybrid map (Fig. 1). A spatial grid map is a top-down grid where each grid cell represents an area in the physical environment. So if an object is at a certain location $(X,Y,Z)$ in a 3D scene, the semantic map will contain information about that object at the corresponding grid cell $(x,y)$, where $x$ and $y$ are the row and column numbers respectively such that there is a direct mapping from $(X,Y,Z)$ to $(x,y)$. For navigation, most spatial maps are 2D, such that a grid cell ignores the $Z$-axis (up direction) by aggregating the semantic information across the up axis. However they can also be 3D where the $Z$-axis is divided into discrete bins. On the other hand, a topological map is a graph-like structure where nodes represent objects or important landmarks in the scene and edges represent relationship (distance, spatial relation, etc.) between them. It is also possible to store semantic information on a dense geometric, which can be viewed as a 3D map with varying density. In a dense geometric, information is associated with each point $(x,y,z)$ corresponding to 3D location $(X,Y,Z)$ in the physical space. Unlike the voxel-grid, which is regularly spaced, points can be sampled at varying densities. Some works combine two or more of the above structures to form a hybrid map since each structure has its own advantages and limitations. We discuss each of these in detail in Sec. 4.
@@ 3.4 How is the map built?
Creating accurate and detailed semantic maps requires integrating data from various sources and sensors such as camera, LiDAR and depth sensors. More specifically, map building consists of having an agent navigate about a space, and accumulating observations $O_{t}$ at time step $t$ into the appropriate map structure $m_{t}$. To build an accurate map, the agent first needs to have an estimate of where it is (localization). Next it extracts semantic information from an observation $F(O_{t})$ (feature extraction), and combines the features into a common map over time (accumulation) (refer to Fig. 5). While building a spatial grid map, an additional step is to project the features onto the map (projection). It is common to group the last three steps into map building and study it jointly with localization in Simultaneous Localization and Mapping (SLAM). We discuss SLAM methods briefly in Sec. 2.2.
@@
Feature extraction. Feature extraction is a crucial part of building a semantic map. Ideally these features should be representative of the objects present in the map. We discuss this topic at length in section Sec. 5.
Projection. An important step in building a spatial grid map is taking the 2D observations and project them into 3D. Typically, this relies on having depth information and known camera parameters in order to convert 2D pixel coordinates to 3D world coordinates. To project a particular pixel in the camera frame, first a ray is shot from the camera center through the image pixel $(i,j)$ to the depth $d_{i,j}$ to get a 3D point in the camera coordinate frame. Next the camera coordinates are converted to the world coordinates $(X,Y,Z)$. For a 2D spatial map, the 3D coordinate$(X,Y,Z)$ is mapped to the grid cell indices $x$ and $y$ in the spatial map. The transformation for the standard pinhole camera with known camera pose (3D rotation $R$ and 3D translation $t$) and intrinsics ($K$) can be written as:
@@ 4 Map structure
In this section we will look at various map structures that have been used in prior works (see Tab. 1) in more depth. A semantic map can be structured in various ways: spatial grid map, topological map, dense geometric or a hybrid map. Spatial grid maps are metric maps of the environment such that its dimensions align to that of the environment and they can be structured as either 2D or 3D grids. Topological maps, on the other hand, represent the environment through a set of landmarks represented as nodes and relation between adjacent landmarks represented as edges in the form of a graph. Dense geometric maps are the densest form of 3D maps whose all three dimensions align to the 3D space such that each 3D point in the scene is captured in the map. Two or more of these types of maps are sometimes combined together to form Hybrid maps.
Figure 6: Spatial map building. A spatial map is a grid map of dimensions ($M\times N\times K$) where $M$ and $N$ are spatial dimensions and $K$ is the number of semantic channels. The common pipeline to build the map is to segment the input image, then ground project into an egocentric map $ego_{map}$, which is then registered to the allocentric map $allo_{map}$ via map update using the localized agent pose. A denoising step generally follows and the map is built online along with the task planning and control. While Gupta et al. (2017) learns a spatial representation without segmenting and ground projecting, Cartillier et al. (2021) observes that encoding followed by ground projecting and then segmenting reduces noise in the produced map.
@@ 4.1 Spatial grid map
A spatial grid map $m_{t}$ is a ($M\times N\times K$) matrix where $M$ and $N$ are the spatial dimensions of the map and $K$ denotes the number of channels to store semantic information at that location. It is a grid like structure where each cell has a width and a height which correspond to a certain area in the physical environment. Current research on indoor embodied AI make use of environments from datasets such as Matterport3D (MP3D) (Chang et al., 2017) and Habitat-Matterport3D (HM3D) (Ramakrishnan et al., 2021) which are 3D reconstructions of real-world spaces. Compared to MP3D, HM3D contains around 10x more scenes with high visual fidelity and lesser reconstruction artifacts. These environments are typically for houses or office spaces with a total area of $<1000\text{m}^{2}$. A grid map with each cell representing an area of $400-900\text{cm}^{2}$ is found to be good enough to represent such spaces (Wani et al., 2020; Raychaudhuri et al., 2023). At the start of each episode, the spatial map is initialized with a tensor of size ($M\times N\times K$), and gradually built as the agent moves around the environment. These are often 2D top-down maps (Gupta et al., 2017; Henriques & Vedaldi, 2018; Narasimhan et al., 2020; Cartillier et al., 2021) with the first two dimensions corresponding to the spatial dimensions of the environment. However, some build 3D spatial maps (Chaplot et al., 2021) to capture the vertical dimension, in which case the map $m_{t}$ is a 4D tensor ($M\times N\times P\times K$).
@@
A spatial map may be built in a number of ways depending on whether raw features are directly projected onto the map, or whether a semantic segmentation is used (see Fig. 6). One way is to learn an egocentric projection of the image features as in CMP (Gupta et al., 2017) which forms the egocentric map. In CMP, the egocentric observations are first encoded with a learned image encoder network such as ResNet (He et al., 2016)), and then the network learns to predict an egocentric projection of the image features, without explicit supervision on the map. Instead, the mapper is trained end-to-end along with the planner to predict actions. Egocentric maps, however, not only fail to capture the global structure of the environment, but ‘forget’ most of the past observations. Thus in long-horizon planning tasks, where the agent needs to ‘remember’ its past observations for efficiency, egocentric maps fall short.
@@
To maintain an allocentric map, it is necessary to take the egocentric information at each time step and aggregate it into a global map. One way to achieve this is to first obtain egocentric projection of image features and then aggregate to a global allocentric map of the environment via a process known as registration. Registration allows the map to incorporate new observations on to specific grid cells. In case the grid cells are already occupied, the new observations are accumulated with the existing ones by employing an aggregation function. This aggregation function can be as simple as taking the latest or the average, but can also be a learned network (see Sec. 3.4). MapNet (Henriques & Vedaldi, 2018) builds an allocentric map by projecting the egocentric image features using depth observations and known camera intrinsics on a 2D top-down grid. This ground projection results in an egocentric projection on a spatial neighborhood around the camera. Next it performs registration by first obtaining a stack of egocentric maps rotated $r$ times and then performing a dense matching with the allocentric map from the previous step to obtain the agent’s current pose on the map. The dense matching is efficiently implemented with convolution operators. A LSTM then performs the aggregation of the current observations rotated by the current pose with the allocentric map from the previous step. While an LSTM is used in this work for aggregation, other functions and neural architectures can be used for aggregation as well (see Tab. 3 for a summary of aggregation methods used in different works).
@@ 4.2 Topological map
Compared to the high-precision grid maps, topological maps are graph-like structures with nodes connected to each other by edges. This essentially abstracts a large space into significant areas (nodes) where the agent can take decisions and connections or paths between them (edges) (Johnson, 2018). This enables parsing the environment into a local and a global structure such that the agent can plan locally in the small space represented as nodes while navigating the large space through graph search following the edges. This way of planning and navigating is inspired from how humans navigate in an unseen environment in that they identify and memorize significant landmarks and find paths to reach those landmarks (Janzen & Van Turennout, 2004; Foo et al., 2005; Chan et al., 2012; Epstein & Vass, 2014). Moreover, in instruction-following tasks, where a robot follows instructions (e.g., ‘go to the yellow chair in front of you’), the language instruction is often represented as a graph with high initial uncertainty. As the robot moves and observes its environment, it uses a probabilistic model to update this graph, leading to improved task performance (Tucker et al., 2019; Patki et al., 2019; Arkin et al., 2020a; Raychaudhuri et al., 2025). Thus topological maps have been a popular choice in traditional robotics research such as SLAM (Thrun & Montemerlo, 2006; Rosinol et al., 2020b; Campos et al., 2021), manipulation (Arkin et al., 2020b; Patki et al., 2020), object search (Aydemir et al., 2011; 2013; Lorbach et al., 2014), language grounding (MacMahon et al., 2006; Chen & Mooney, 2011b; Tellex et al., 2011a; Matuszek et al., 2013; Duvallet et al., 2013; Gong & Zhang, 2018; Paul et al., 2018) as well as in embodied AI research (Savinov et al., 2018; Chen et al., 2021; Chaplot et al., 2020b; Kwon et al., 2021; Gu et al., 2023b; Mehan et al., 2024; Garg et al., 2024; An et al., 2024; Yang et al., 2024; Tang et al., 2025).
@@
A key design decision during topological map building is what should be represented as nodes and what should be edges. Generally speaking, the nodes encode semantic information about locations in the environment such that the agent can make a decision whereas the edges store relationship or connection between the nodes. For indoor navigation, the landmarks for the nodes are typically objects in the environments. They can also be openings or intersections (Fredriksson et al., 2023), locations the agents has visited (Chaplot et al., 2020b), and other regions of interest (Kim et al., 2023; Shah et al., 2023; Garg et al., 2024). For navigation, two nodes are connected with an edge if it is possible to navigate from one node to another. Some methods also store spatial relationships between the nodes (Gu et al., 2023b) in the edges to enable better reasoning.
One way to construct a topological map (see Fig. 6) is during an exploration phase previous to the actual task and then use the graph to plan a path to the node most similar to the target, for example, in Semi-Parametric Topological Memory (SPTM) (Savinov et al., 2018). During exploration the agent follows multiple random trajectories for each environment to form a node for every visited location and add an edge between the current node and the previous one to encode connectivity or reachability between them. A common post-processing step includes trimming out redundant nodes and edges to form a sparse graph (Chen et al., 2021). When the graph of one environment is collected from multiple random trajectories, it is also common to merge these graphs into one. However, a topological map generated this way in a pre-exploration phase is still sparse meaning that some observations in the environment might not have been captured by the graph. This affects the agent performance in the downstream task. Moreover, they need a pre-exploration phase which makes them unsuitable for unseen environments.
@@
Another important aspect in the topological map creation is how to determine if two observations are similar to each other, in which case the two are mapped to the same node. If they are not similar, two different nodes exist for the two observations. This requires the map building methods to compare RGB images. The goal here is to classify two images as similar if (1) they are exactly the same or (2) there is a slight change in direction or distance between the two. Traditionally this is the problem of data association in SLAM-based systems, where an incoming observation could be matched to multiple landmarks (nodes) and either the best match is selected or a new node is created to mark a new landmark (Bowman et al., 2017; Dellaert et al., 2000). In embodied AI some works use a pretrained classifier network to implement this. The network is trained to classify whether two images are from the same area. NTS uses MLP trained with a cross-entropy loss in a supervised manner to predict whether are similar. This however needs annotated pairs of training data. Cross-Modal Transformer Planner (CMTP) (Chen et al., 2021), on the other hand, uses an oracle ‘Reachability Estimator’ to first obtain the geodesic distance between the two underlying locations based on the traversibility of the 3D mesh. If the distance is below a threshold, it maps them to the same node. Visual Graph Memory (VGM) (Kwon et al., 2021) also uses a pretrained network to determine if two images are similar. But they learn an unsupervised representation of the observations which are then projected onto an embedding space. The idea is to have the embeddings of observations coming from nearby areas clustered together because they are likely to have similar appearances. The training data in this case consist of randomly sampled observations from the training environments, thus eliminating the need for manual annotations. Kim et al. (2023) on the other hand use semantic similarity score obtained from a pretrained network (Li et al., 2021a) between two images to determine whether they are the same nodes. A similar approach is taken in LM-Nav (Shah et al., 2023) where they use CLIP to calculate the cosine similarity between image features. Tab. 4 compares different methods that build topological maps.

### v3

@@ 2.3.2 Modular pipeline
Another line of work explores how to breakdown a complex navigation task into a set of basic skills that the agent needs to acquire. Such skills can then be learned independently of each other so that they can be leveraged across various tasks without the need to be retrained from scratch. While some use two modules (Bansal et al., 2020), one for high-level decision making and another for low-level planner, others (Chaplot et al., 2019; 2020a; Gervet et al., 2023; Raychaudhuri et al., 2023) further breakdown into sub-modules, such as visual encoder, mapper and exploration in addition to the low-level planner. Fig.˜4 shows the difference between the end-to-end approach and the module approach, and we elaborate on the modules and popular design choices below.
Visual Encoder. This module encodes agent observations to produce semantic visual features and predictions at every time step. Prior works have used visual features from pretrained backbones such as ResNet He et al. (2016) or ViT (Dosovitskiy et al., 2020), and often leveraging object detectors MaskRCNN He et al. (2017a) or FasterRCNN Ren et al. (2015). As we will see in Sec.˜5.2.2, with the development of large pretrained vision-language models and open-vocabulary detectors, pretrained models such as CLIP (Radford et al., 2021), LSeg (Li et al., 2022), DINO (Caron et al., 2021; Oquab et al., 2023), and others are increasingly popular as the basis for building large to open-vocabulary maps. The visual encoder used will determine the information captured in the features, and whether there are detected object instance bounding boxes or segmentations for integration into the mapping modules.
Mapper. The mapper is responsible for building a semantic map of the environment from the encoded image features and agent pose. To build a global map over time, the mapper typically aggregates the current map with the map from previous step (see Sec.˜3.4 for details). In this paper, we survey how recent methods structure the map (Sec.˜4) and what information can be encoded in it (Sec.˜5). Tab.˜2 summarizes the type of information stored in the map by various methods. This can be occupancy information, explored area or semantic labels of the detected objects.
@@
In Tab.˜2, we compare common modular approaches and look at how they leverage various heuristics-based or learned approach for each of the modules. The advantages of a modular pipeline include its ability to leverage pretrained models from other tasks (Gervet et al., 2023; Raychaudhuri et al., 2023) and its ability to transfer from simulation to real-world robots better (Gervet et al., 2023).
@@ 3 Semantic map
In this section, we introduce semantic maps (Sec.˜3.1), and provide an overview of their structural representations (Sec.˜3.2), semantic encodings (Sec.˜3.3), and the process of building them (Sec.˜3.4).
@@ 3.1 What are semantic maps?
While maps, capturing the geometry of the space, can help an agent avoid obstacles, they are not sufficient for the demands of more complex reasoning tasks. An enhanced map that goes beyond geometry to capture meaning and context in its environment aligns with how humans perceive and navigate its surroundings. These are semantic maps, which provide a richer and more nuanced understanding about the objects and places in the environment and are indispensable for performing complex tasks such as navigating to a specific room (kitchen) (Narasimhan et al., 2020), rearranging objects (Trabucco et al., 2022) or performing an action on a specific object (sitting on a couch) (Peng et al., 2023). A physical or a virtual agent perceives the environment through sensors (camera, LiDAR, etc.), uses cognition to classify the objects and regions it perceives (Ren et al., 2015; He et al., 2017a; Liu et al., 2024; Kirillov et al., 2023; Zhang et al., 2023b), stores them in a structured semantic map which can be queried for efficient reasoning.
@@ 3.2 What is the structure of this map?
A semantic map can be structured as a spatial grid map, topological map, dense geometric or a hybrid map (Fig.˜1). A spatial grid map is a top-down grid where each grid cell represents an area in the physical environment. So if an object is at a certain location $(X,Y,Z)$ in a 3D scene, the semantic map will contain information about that object at the corresponding grid cell $(x,y)$, where $x$ and $y$ are the row and column numbers respectively such that there is a direct mapping from $(X,Y,Z)$ to $(x,y)$. For navigation, most spatial maps are 2D, such that a grid cell ignores the $Z$-axis (up direction) by aggregating the semantic information across the up axis. However they can also be 3D where the $Z$-axis is divided into discrete bins. On the other hand, a topological map is a graph-like structure where nodes represent objects or important landmarks in the scene and edges represent relationship (distance, spatial relation, etc.) between them. It is also possible to store semantic information on a dense geometric, which can be viewed as a 3D map with varying density. In a dense geometric, information is associated with each point $(x,y,z)$ corresponding to 3D location $(X,Y,Z)$ in the physical space. Unlike the voxel-grid, which is regularly spaced, points can be sampled at varying densities. Some works combine two or more of the above structures to form a hybrid map since each structure has its own advantages and limitations. We discuss each of these in detail in Sec.˜4.
@@ 3.4 How is the map built?
Creating accurate and detailed semantic maps requires integrating data from various sources and sensors such as camera, LiDAR and depth sensors. More specifically, map building consists of having an agent navigate about a space, and accumulating observations $O_{t}$ at time step $t$ into the appropriate map structure $m_{t}$. To build an accurate map, the agent first needs to have an estimate of where it is (localization). Next it extracts semantic information from an observation $F(O_{t})$ (feature extraction), and combines the features into a common map over time (accumulation) (refer to Fig.˜5). While building a spatial grid map, an additional step is to project the features onto the map (projection). It is common to group the last three steps into map building and study it jointly with localization in Simultaneous Localization and Mapping (SLAM). We discuss SLAM methods briefly in Sec.˜2.2.
@@
Feature extraction. Feature extraction is a crucial part of building a semantic map. Ideally these features should be representative of the objects present in the map. We discuss this topic at length in section Sec.˜5.
Projection. An important step in building a spatial grid map is taking the 2D observations and project them into 3D. Typically, this relies on having depth information and known camera parameters in order to convert 2D pixel coordinates to 3D world coordinates. To project a particular pixel in the camera frame, first a ray is shot from the camera center through the image pixel $(i,j)$ to the depth $d_{i,j}$ to get a 3D point in the camera coordinate frame. Next the camera coordinates are converted to the world coordinates $(X,Y,Z)$. For a 2D spatial map, the 3D coordinate(X,Y,Z)(X,Y,Z) is mapped to the grid cell indices $x$ and $y$ in the spatial map. The transformation for the standard pinhole camera with known camera pose (3D rotation $R$ and 3D translation $t$) and intrinsics ($K$) can be written as:
@@
Figure 6: This timeline highlights how semantic mapping methods have progressed over the years, with increasing diversity in map structure and encoding. While various map structures have long been explored, the last few years have seen a shift toward open-vocabulary semantic maps, driven by advances in large language and vision-language models.
@@ 4 Map structure
In this section we will look at various map structures that have been used in prior works (see Tab.˜1) in more depth. A semantic map can be structured in various ways: spatial grid map, topological map, dense geometric or a hybrid map. Spatial grid maps are metric maps of the environment such that its dimensions align to that of the environment and they can be structured as either 2D or 3D grids. Topological maps, on the other hand, represent the environment through a set of landmarks represented as nodes and relation between adjacent landmarks represented as edges in the form of a graph. Dense geometric maps are the densest form of 3D maps whose all three dimensions align to the 3D space such that each 3D point in the scene is captured in the map. Two or more of these types of maps are sometimes combined together to form Hybrid maps.
Figure 7: Grid map building. A spatial grid map has dimensions ($M\times N\times K$) where $M$ and $N$ are spatial dimensions and $K$ is the number of semantic channels. The common pipeline to build the map is to segment the input image, then ground project into an egocentric map $ego_{map}$, which is then registered to the allocentric map $allo_{map}$ via map update using the localized agent pose. A denoising step generally follows and the map is built online along with the task planning and control. While Gupta et al. (2017) learns a spatial representation without segmenting and ground projecting, Cartillier et al. (2021) observes that encoding followed by ground projecting and then segmenting reduces noise in the produced map.
@@
Dense geometric map building. A dense geometric map accumulates semantic information directly onto the 3D geometry of a scene, most popularly using a point-cloud.
@@ 4.1 Spatial grid map
A spatial grid map $m_{t}$ is a ($M\times N\times K$) matrix where $M$ and $N$ are the spatial dimensions of the map and $K$ denotes the number of channels to store semantic information at that location. It is a grid like structure where each cell has a width and a height which correspond to a certain area in the physical environment. Current research on indoor embodied AI make use of environments from datasets such as Matterport3D (MP3D) (Chang et al., 2017) and Habitat-Matterport3D (HM3D) (Ramakrishnan et al., 2021) which are 3D reconstructions of real-world spaces. Compared to MP3D, HM3D contains around 10x more scenes with high visual fidelity and lesser reconstruction artifacts. These environments are typically for houses or office spaces with a total area of $1000\text{m}^{2}$. A grid map with each cell representing an area of $400-900\text{cm}^{2}$ is found to be good enough to represent such spaces (Wani et al., 2020; Raychaudhuri et al., 2023). At the start of each episode, the spatial map is initialized with a tensor of size ($M\times N\times K$), and gradually built as the agent moves around the environment. These are often 2D top-down maps (Gupta et al., 2017; Henriques & Vedaldi, 2018; Narasimhan et al., 2020; Cartillier et al., 2021) with the first two dimensions corresponding to the spatial dimensions of the environment. However, some build 3D spatial maps (Chaplot et al., 2021) to capture the vertical dimension, in which case the map $m_{t}$ is a 4D tensor ($M\times N\times P\times K$).
@@
A spatial map may be built in a number of ways depending on whether raw features are directly projected onto the map, or whether a semantic segmentation is used (see Fig.˜7). One way is to learn an egocentric projection of the image features as in CMP (Gupta et al., 2017) which forms the egocentric map. In CMP, the egocentric observations are first encoded with a learned image encoder network such as ResNet (He et al., 2016)), and then the network learns to predict an egocentric projection of the image features, without explicit supervision on the map. Instead, the mapper is trained end-to-end along with the planner to predict actions. Egocentric maps, however, not only fail to capture the global structure of the environment, but ‘forget’ most of the past observations. Thus in long-horizon planning tasks, where the agent needs to ‘remember’ its past observations for efficiency, egocentric maps fall short.
@@
To maintain an allocentric map, it is necessary to take the egocentric information at each time step and aggregate it into a global map. One way to achieve this is to first obtain egocentric projection of image features and then aggregate to a global allocentric map of the environment via a process known as registration. Registration allows the map to incorporate new observations on to specific grid cells. In case the grid cells are already occupied, the new observations are accumulated with the existing ones by employing an aggregation function. This aggregation function can be as simple as taking the latest or the average, but can also be a learned network (see Sec.˜3.4). MapNet (Henriques & Vedaldi, 2018) builds an allocentric map by projecting the egocentric image features using depth observations and known camera intrinsics on a 2D top-down grid. This ground projection results in an egocentric projection on a spatial neighborhood around the camera. Next it performs registration by first obtaining a stack of egocentric maps rotated $r$ times and then performing a dense matching with the allocentric map from the previous step to obtain the agent’s current pose on the map. The dense matching is efficiently implemented with convolution operators. A LSTM then performs the aggregation of the current observations rotated by the current pose with the allocentric map from the previous step. While an LSTM is used in this work for aggregation, other functions and neural architectures can be used for aggregation as well (see Tab.˜3 for a summary of aggregation methods used in different works).
@@ 4.2 Topological map
Compared to the high-precision grid maps, topological maps are graph-like structures with nodes connected to each other by edges. This essentially abstracts a large space into significant areas (nodes) where the agent can take decisions and connections or paths between them (edges) (Johnson, 2018). This enables parsing the environment into a local and a global structure such that the agent can plan locally in the small space represented as nodes while navigating the large space through graph search following the edges. This way of planning and navigating is inspired from how humans navigate in an unseen environment in that they identify and memorize significant landmarks and find paths to reach those landmarks (Janzen & Van Turennout, 2004; Foo et al., 2005; Chan et al., 2012; Epstein & Vass, 2014). Moreover, in instruction-following tasks, where a robot follows instructions (e.g., ‘go to the yellow chair in front of you’), the language instruction is often represented as a graph with high initial uncertainty. As the robot moves and observes its environment, it uses a probabilistic model to update this graph, leading to improved task performance (Tucker et al., 2019; Patki et al., 2019; Arkin et al., 2020a; Raychaudhuri et al., 2025). Thus topological maps have been a popular choice in traditional robotics research such as SLAM (Thrun & Montemerlo, 2006; Rosinol et al., 2020b; Campos et al., 2021), manipulation (Arkin et al., 2020b; Patki et al., 2020), object search (Aydemir et al., 2011; 2013; Lorbach et al., 2014), language grounding (MacMahon et al., 2006; Chen & Mooney, 2011; Tellex et al., 2011a; Matuszek et al., 2013; Duvallet et al., 2013; Gong & Zhang, 2018; Paul et al., 2018) as well as in embodied AI research (Savinov et al., 2018; Chen et al., 2021; Chaplot et al., 2020b; Kwon et al., 2021; Gu et al., 2024; Mehan et al., 2024; Garg et al., 2024; An et al., 2024; Yang et al., 2024; Tang et al., 2025).
@@
A key design decision during topological map building is what should be represented as nodes and what should be edges. Generally speaking, the nodes encode semantic information about locations in the environment such that the agent can make a decision whereas the edges store relationship or connection between the nodes. For indoor navigation, the landmarks for the nodes are typically objects in the environments. They can also be openings or intersections (Fredriksson et al., 2023), locations the agents has visited (Chaplot et al., 2020b), and other regions of interest (Kim et al., 2023; Shah et al., 2023; Garg et al., 2024). For navigation, two nodes are connected with an edge if it is possible to navigate from one node to another. Some methods also store spatial relationships between the nodes (Gu et al., 2024) in the edges to enable better reasoning.
One way to construct a topological map (see Fig.˜7) is during an exploration phase previous to the actual task and then use the graph to plan a path to the node most similar to the target, for example, in Semi-Parametric Topological Memory (SPTM) (Savinov et al., 2018). During exploration the agent follows multiple random trajectories for each environment to form a node for every visited location and add an edge between the current node and the previous one to encode connectivity or reachability between them. A common post-processing step includes trimming out redundant nodes and edges to form a sparse graph (Chen et al., 2021). When the graph of one environment is collected from multiple random trajectories, it is also common to merge these graphs into one. However, a topological map generated this way in a pre-exploration phase is still sparse meaning that some observations in the environment might not have been captured by the graph. This affects the agent performance in the downstream task. Moreover, they need a pre-exploration phase which makes them unsuitable for unseen environments.
@@
Another important aspect in the topological map creation is how to determine if two observations are similar to each other, in which case the two are mapped to the same node. If they are not similar, two different nodes exist for the two observations. This requires the map building methods to compare RGB images. The goal here is to classify two images as similar if (1) they are exactly the same or (2) there is a slight change in direction or distance between the two. Traditionally this is the problem of data association in SLAM-based systems, where an incoming observation could be matched to multiple landmarks (nodes) and either the best match is selected or a new node is created to mark a new landmark (Bowman et al., 2017; Dellaert et al., 2000). In embodied AI some works use a pretrained classifier network to implement this. The network is trained to classify whether two images are from the same area. NTS uses MLP trained with a cross-entropy loss in a supervised manner to predict whether are similar. This however needs annotated pairs of training data. Cross-Modal Transformer Planner (CMTP) (Chen et al., 2021), on the other hand, uses an oracle ‘Reachability Estimator’ to first obtain the geodesic distance between the two underlying locations based on the traversibility of the 3D mesh. If the distance is below a threshold, it maps them to the same node. Visual Graph Memory (VGM) (Kwon et al., 2021) also uses a pretrained network to determine if two images are similar. But they learn an unsupervised representation of the observations which are then projected onto an embedding space. The idea is to have the embeddings of observations coming from nearby areas clustered together because they are likely to have similar appearances. The training data in this case consist of randomly sampled observations from the training environments, thus eliminating the need for manual annotations. Kim et al. (2023) on the other hand use semantic similarity score obtained from a pretrained network (Li et al., 2021a) between two images to determine whether they are the same nodes. A similar approach is taken in LM-Nav (Shah et al., 2023) where they use CLIP to calculate the cosine similarity between image features. Tab.˜4 compares different methods that build topological maps.